# On the Need for a Statistical Foundation in Scenario-Based Testing of Autonomous Vehicles

### v1

@@ Abstract
Scenario-based testing has emerged as a common method for autonomous vehicles (AVs) safety, offering a more efficient alternative to mile-based testing by focusing on high-risk scenarios. However, fundamental questions persist regarding its stopping rules, residual risk estimation, debug effectiveness, and the impact of simulation fidelity on safety claims. This paper argues that a rigorous statistical foundation is essential to address these challenges and enable rigorous safety assurance. By drawing parallels between AV testing and traditional software testing methodologies, we identify shared research gaps and reusable solutions. We propose proof-of-concept models to quantify the probability of failure per scenario (pfs) and evaluate testing effectiveness under varying conditions. Our analysis reveals that neither scenario-based nor mile-based testing universally outperforms the other. Furthermore, we introduce Risk Estimation Fidelity (REF), a novel metric to certify the alignment of synthetic and real-world testing outcomes, ensuring simulation-based safety claims are statistically defensible.
@@ I Introduction
Autonomous vehicles (AVs) are transitioning rapidly from research environments to public roads. Waymo initiated this shift by launching its first commercial AV taxi in December 2018. More recently, Tesla has announced plans to deploy a robotaxi service in Austin, Texas, by June 2025, leveraging its Full Self-Driving system. These developments underscore the urgent need for comprehensive and rigorous safety assessment to satisfy regulatory standards and public safety expectations.
While many efforts have been made to address the AV safety challenges covering various aspects like design, implementation, regulation and legal issues [1, 2, 3, 4, 5], AV testing remains an indispensable method for generating the verification and validation evidence necessary for safety assurance. In general, there are two common types of AV testing, mile-based and scenario-based. The former involves accumulating a large number of real-world/simulated miles to statistically validate the vehicle’s safety. The premise is that, as the vehicle encounters a wider variety of driving conditions and edge cases, statistical confidence in its safety increases. Given the required rarity of incidents, demonstrating safety through this method would require millions/billions of miles [6, 7, 8]. The industry largely resists such mile-based “driving to safety” strategy due to the prohibitive cost and risk. Instead, they favour scenario-based testing [9, 10, 11, 12] as a complementing approach, which focuses on using controlled, high-risk scenarios to infer the AV’s safety. This method avoids the need for extensive driven miles; instead, it tests the vehicle’s response to predefined, potentially dangerous situations through simulated or field-tested extreme scenarios.
@@
i) What is the stopping rule of scenario-based testing? The number of possible parameter combinations to describe test scenarios is exploding, so exhaustive testing of all possible scenarios for a given operational design domain (ODD) is infeasible [13]. While coverage criteria were studied (e.g. in [13, 14, 15]) for scenario-based testing, borrowing ideas from traditional software testing where coverage criteria are used to decide when to stop [16], problems like the arbitrariness of discretising continuous parameters and redundancy/bias in scenario generation limit their effectiveness [13].
@@
ii) What is the residual safety risk after extensive scenario-based testing? Even after testing a vast number of scenarios (say achieved 100% coverage of some criteria), the residual risk (measured by, e.g., the probability of crash per random scenario/mile) remains unknown. This uncertainty arises because coverage criteria typically measure which parts of the input space have been exercised but do not translate directly into an estimate of overall system safety delivered  [17, 18] (a well-known problem in software engineering).
@@
iii) Is scenario-based testing always more effective at improving safety of AVs? Scenario-based testing “smartly” selects conditions that are hypothesised to stress the AV, thus potentially detecting more bugs than mile-based testing. However, “more bugs does not necessarily imply less reliable” [19]. Consider an extreme example: if scenario-based testing uncovers only “5,000-year bugs” [^1] E. Adams studied fault occurrence rates in large IBM software systems and found that over 60% were “5,000-year bugs”, meaning each fault manifests as failures, on average, once every 5,000 years in normal operation [20]., fixing them may have little improvement on AVs’ operational safety, compared to fixing fewer “5-year bugs” discovered through mile-based testing. Essentially, not all bugs contribute equally to operational safety risk. Scenario-based testing, e.g., using optimisation algorithms to achieve a required level of coverage, often overlooks this fact in its objectives, while mile-based testing naturally detects more “common bugs”.
@@
iv) How does the fidelity [^2] While specific definitions of fidelity vary across studies, in general, fidelity refers to how closely synthetic data replicates real-world conditions, thereby supporting reliable simulation-based testing of AVs. of simulated scenarios affect safety estimates? Given that scenario-based testing focuses on high-risk conditions, synthetic data generated by simulation is typically employed to execute test scenarios. However, the Sim2Real gap introduces an additional layer of uncertainty, implying that safety claims based on synthetic scenario testing evidence may not fully capture real-world performance [10]. Despite great efforts having been made to bridge the gap [21, 22, 23, 24], there remains a lack of rigorous methods to quantify simulation fidelity [23] and translate varying fidelity levels into measurable safety risk.
@@
In this paper, we begin by reviewing the state-of-the-art in statistical modelling for AV safety assessment. We observe that most existing work focuses on mile-based testing, while statistical foundations for scenario-based testing remain largely under-explored, despite its growing prominence in industry practice. Furthermore, we notice that many of the open questions in scenario-based testing, e.g., effectiveness in debugging and residual risk estimation, closely parallel longstanding challenges studied in the traditional software testing community. We present an initial model, as a proof-of-concept, to demonstrate the potential of applying statistical approaches to rigorously reason about scenario-based safety evidence, and to highlight the potential for AV testing to benefit from software testing, with existing models potentially being reusable. Our goal is to lay a foundation that future work can build upon to develop rigorous safety arguments for AVs.
@@
  • Proof-of-concept models using scenario-based testing evidence to assess AV safety and evaluate their effectiveness in reducing operational risk.
  • A novel notion of Risk Estimation Fidelity of synthetic data, with formal definitions and example use cases.
@@ II-A AV Testing
Mile-based testing evaluates AV safety by accumulating real/simulated miles within the ODD, aiming to sample scenarios as they naturally occur. Safety is inferred statistically by analysing failure rates (e.g., crashes, disengagements) per unit distance. It assumes an unbiased estimate and exposure to diverse scenarios, potentially offering some benefits similar to scenario-based testing.
@@ II-B Software Testing
Broadly, software testing falls into two categories: directed testing and operational testing; they are also often called debug testing and acceptance testing [18, 26] based on their intended purposes. The former aims to find and fix defects, while the latter assesses existing quality to build confidence for deployment. The key difference is whether test cases are sampled from the operational profile (OP). Debug testing typically ignores the OP, whereas acceptance testing depends on it [18]. However, research shows operational testing may outperform directed testing in bug detection, and some acceptance tests rely solely on directed methods. We show this using our proof-of-concept models later.
@@
An OP is a probability distribution over the input space of a software system, representing the relative frequencies with which different inputs are expected to occur during actual operation.
@@
Partition-based testing partitions the input domain of a software system into distinct subsets. Test cases are then selected from each partition to ensure comprehensive coverage of the input space.
@@ II-C Mapping Concepts Between AV and Software Testing
Scenario-based testing can be viewed as a form of partition-based testing, where logical scenarios partition the ODD, and representative concrete scenarios are tested in each partition to ensure coverage. Both aim to save tests by assuming failures cluster in certain “smartly partitioned” subdomains, allowing testers to focus effort where it matters most.
Remark 4 (Mile-based testing vs random testing).
Continuing Remark 1, since each concrete scenario arises randomly from an operational scenario distribution (i.e., the OP), mile-based testing resembles repeated bursts of random testing across the scenario parameter space.
@@
That said, there are exceptional cases that we learnt from studies of software testing, e.g., [18, 19, 8, 17, 26, 28, 29]: 1) Partition/scenario-based testing is often thought-to-be good at debugging, but there is no formal proof and it may be less effective under certain, not especially unusual, conditions. 2) Partition/scenario-based testing can still be used for estimating reliability (i.e., residual safety risk), if distributional information is explicitly incorporated within and across logical scenario partitions. 3) Random/mile-based testing can still discover bugs, whether they may or may not be less efficient in improving safety. All of these points are context-dependent and cannot be answered universally. This highlights the need for a theoretical model with statistical foundation capable of analysing different AV testing methods across conditions.
@@ III Statistical Modelling for AV Safety
[7] [36] pfm Bernoulli CBI Failure-free & rare failure illustrative data In addition to [6], prior knowledge may bring advantages, while avoiding optimistic biases.
@@
[39] Bounded reachability probability Poisson Model checker on SAN HighD dataset [40], Lyft [41] Probabilistic framework to understand and mitigate AV safety risks.
@@
The nature of the metric, whether it represents a probability, e.g., probability of failure per mile (pfm), or a rate, e.g., disengagements per mile (dpm), is crucial in determining the appropriate modelling approach. Although probability and rate are sometimes used interchangeably, they are fundamentally different. Probability represents the likelihood of an event occurring and ranges from 0 to 1. In contrast, rate quantifies the number of events per selected unit (e.g., time or distance), and ranges from 0 to $\infty$. It determines the choice of the underlying stochastic process (e.g., Bernoulli for probabilities or Poisson for rates) and informs whether a discrete or continuous model is more appropriate. The metric also guides how data should be collected in practice, and whether the data realistically supports the assumptions of the chosen stochastic process.
@@
A Stochastic Activity Network (SAN) is a probabilistic modelling framework which is used to represent complex, stochastic systems. It can be used to assess the safety of AVs under various hazardous conditions, e.g., road hazards, cyber-attacks, etc. There are tools (like Möbius) that help analyse and simulate SAN models.
@@
Frequentist methods, such as statistical significance tests, rely only on observed data, which can make them less reliable when dealing with small datasets. In the case of AVs, this can be challenging as a large amount of data is often needed for safety demonstration. They also cannot incorporate prior knowledge, an advantage offered by Bayesian inference methods (e.g., CBI). However, Bayesian estimators are normally sensitive to prior knowledge and how it is translated into mathematical form [46].
@@
SRGMs help track and predict failures over time, but they cannot be trusted a priori. They are especially problematic for highly reliable software and, even when generally accurate for a system, can occasionally produce seriously incorrect predictions. For SAN models, accurately mapping real-world components and data into SAN structures and transition rates between states is crucial and challenging.
@@ IV-A The Metric, Stochastic Failure Processes, and Estimators
As stated in Remark 6, the choice of metric is crucial in statistical inference. As a first attempt, we align with existing metrics and define the probability of failure per randomly selected scenario (drawn from the operational scenario distribution over the ODD), abbreviated as pfs:
@@
Similar to the “frequentist” interpretation of pfd in [47], pfs can be regarded as a limiting relative frequency of concrete scenarios for which the AV fails in an infinite sequence of independently selected concrete scenarios from the OP. Note, as stated in footnote 3, the “failure” in pfs is a general term that can be replaced by any safety event of the assessor’s interest, e.g., crash and near-miss.
@@ IV-A1 Scenario-based testing conducted in a mile-based/random testing way
Given the likelihood function, many statistical inference models can be better used for estimating $\theta$, including those summarised in Table I, depending on the observed data (e.g., failure-free), prior knowledge available, and desired posterior estimation forms (e.g., mean or confidence bounds of pfs). E.g., a Bayesian estimator for the posterior mean of pfs is:
@@ IV-A2 Scenario-based testing conducted in a partition-based testing way
Let the scenario parameter space $D$ partitioned into subdomains $D_{1},\dots,D_{n}$, i.e., $n$ logical scenarios. Then, by the law of total probability
@@
Thus, to estimate $\theta$, we need to know the conditional pfs of each logical scenario $\theta_{i}$ and the OP. Normally, inside a given logical scenario $D_{i}$, a new and (thought-to-be) more efficient distribution is used to generate concrete test scenarios (e.g., by an optimisation algorithm) rather than the (local) OP. Usually, the estimator using importance sampling can be used where the proposal distribution is either known or implicitly approximated [30, 31, 32, 33].
@@ IV-B Under Which Conditions Will Scenario-based Testing Deliver Better Safety than Mile-based Testing?
We continue to investigate the above question by retrofitting the model from [18] for AV testing. To model the fix behaviour after detecting failures, [18] introduces the concept of “failure regions ($F$)”—sets of inputs in $D$ that cause failure and are eliminated by fixes. It assumes [^5] The original paper [18] discusses the validity of some assumptions and model extensions to relax them, e.g., “partial fix”. For simplicity, we omit these discussions here. a fix removes a specific failure region and does not introduce new failure regions. While [18] considers the more general case of multiple (disjoint) failure regions, we only present the simplified single failure region case to better illustrate the conceptual mapping between AV and software testing, and the potential for reusing existing research insights.
Consider an AV under testing with pfs, $\theta=q$. After doing mile-based testing that included $t$ random concrete scenarios (cf. Remarks 1 & 4). Depending on whether the failure region has been detected or not, the expected pfs after fixing is stated in Eq. (5), where the “00” reflects the fact that after fixing, there is no bug residing.
@@
Consider the failure region is a strict subset of a single subdomain $k$, $F\subset D_{k}$, and with some points in $D_{k}$ not being failure points, $D_{k}\not\subset F$, and no other subdomains overlapping with $F$: $F\cap D_{i}=\emptyset,\forall i\neq k$. Additionally, we assume that within $D_{k}$, both scenario-based and mile-based testing (on average) have an equal probability to encounter $F$. That is, the detection probability $d_{k}$ is just the fraction of the operational distribution scenarios in $D_{k}$ that encounter $F$: $d_{k}=\frac{\sum_{x\in F}Op(x)}{\sum_{x\in D_{k}}Op(x)},$ For scenario-based testing, assume test budgets are distributed equally across $n$ subdomains, so $t_{k}=t/n$, where $t$ is the total number of tests.
  • Situations when scenario-based testing is superior: This case occurs when the failure-prone subdomain is relatively rare in real operation, so that the mile-based testing gives it fewer tests than the scenario-based testing. Mathematically, this is expressed as $t\sum_{x\in D_{k}}Op(x)\ll t_{k}$. After substituting $t_{k}=t/n$, we will have:
@@
    where the last term $q(1-q)^{t}$ is the pfs for mile-based testing. The two approximations in Eq.s (12) require that $d_{k}$ and $q$ are small, using $(1+x)^{y}\approx 1+yx$ for small $x$.
@@
  • Situations when mile-based testing is superior: A similar argument leads to the opposite conclusion when mile-based testing heavily samples $D_{k}$ (many mile-based test scenarios fall within $D_{k}$). If there are many other subdomains, scenario-based testing ends up “wasting” most of its effort on them (still assuming $t_{k}=t/n$), rather than the “failure prone” subdomain. Mathematically, mile-based testing provides more tests, $t\sum_{x\in D_{k}}Op(x)\gg t_{k}$, substituting $t_{k}=t/n$, we have:

### v2

@@ Abstract
Scenario-based testing has emerged as a common method for autonomous vehicles (AVs) safety assessment, offering a more efficient alternative to mile-based testing by focusing on high-risk scenarios. However, fundamental questions persist regarding its stopping rules, residual risk estimation, debug effectiveness, and the impact of simulation fidelity on safety claims. This paper argues that a rigorous statistical foundation is essential to address these challenges and enable rigorous safety assurance. By drawing parallels between AV testing and established software testing methods, we identify shared research gaps and reusable solutions. We propose proof-of-concept models to quantify the probability of failure per scenario (pfs) and evaluate testing effectiveness under varying conditions. Our analysis reveals that neither scenario-based nor mile-based testing universally outperforms the other. Furthermore, we give an example of formal reasoning about alignment of synthetic and real-world testing outcomes, a first step towards supporting statistically defensible simulation-based safety claims.
@@
{textblock*}
20cm(1cm,1cm) Accepted by ITSC 2025: https://ieee-itsc.org/2025/
@@ I Introduction
Autonomous vehicles (AVs) are transitioning rapidly from research environments to public roads. Waymo initiated this shift by launching its first commercial AV taxi in December 2018. More recently, Tesla has announced plans to deploy a robotaxi service in Austin, Texas, by June 2025, using its Full Self-Driving system. These developments underscore the urgent need for comprehensive and rigorous safety assessment to satisfy regulatory standards and public safety expectations.
While many efforts have been made to address the AV safety challenges covering various aspects like design, implementation, regulation and legal issues [1, 2, 3, 4, 5], AV testing remains an indispensable method for generating the verification and validation evidence necessary for safety assurance. In general, there are two common types of AV testing, mile-based and scenario-based [^1] Scenario-based testing may be executed by conducting the tests on different “platforms”: simulators, “proving grounds” or on the public roads.. The former involves accumulating a large number of real-world or simulated miles to statistically validate the vehicle’s safety. The premise is that, as the vehicle encounters a wider variety of driving conditions and edge cases, statistical confidence in its safety increases. Given the required rarity of incidents, demonstrating safety through this method would require millions/billions of miles [6, 7, 8]. The industry largely resists such mile-based “driving to safety” strategy due to the prohibitive cost and risk. Instead, they favour scenario-based testing [9, 10, 11, 12] as a complementary approach, which focuses on using controlled, high-risk scenarios to infer the AV’s safety. This method avoids the need for extensive driven miles; instead, it tests the vehicle’s response to predefined, potentially dangerous situations through simulated or field-tested extreme scenarios.
@@
i) What is the stopping rule of scenario-based testing? The number of possible parameter combinations to describe test scenarios is exploding, so exhaustive testing of all possible scenarios for a given operational design domain (ODD) is infeasible [13]. While coverage criteria were studied (e.g. in [13, 14, 15]) for scenario-based testing, borrowing ideas from traditional software testing where coverage criteria are used to decide when to stop [16], problems like the arbitrariness of discretising continuous parameters and redundancy/bias in scenario generation limit their credibility [13].
@@
ii) What is the residual safety risk after extensive scenario-based testing? Even after testing a vast number of scenarios (say achieved $100\%$ coverage of some criteria), the residual risk (measured by, e.g., the probability of crash per random scenario/mile) remains unknown. This uncertainty arises because coverage criteria typically measure which parts of the input space have been exercised but do not translate directly into an estimate of overall system safety delivered  [17, 18] (a well-known problem in software engineering).
@@
iii) Is scenario-based testing always more effective at improving safety of AVs? Scenario-based testing “smartly” selects conditions that are hypothesised to stress the AV, thus potentially detecting more bugs than mile-based testing. However, “more bugs does not necessarily imply less reliable” [19]. Consider an extreme example: if scenario-based testing uncovers only “5,000-year bugs” [^2] E. Adams studied fault occurrence rates in large IBM software systems and found that over 60% were “5,000-year bugs”, meaning each fault manifests as failures, on average, once every 5,000 years of operation [20]., fixing them may have little improvement on AVs’ operational safety, compared to fixing fewer “5-year bugs” discovered through mile-based testing. Essentially, not all bugs contribute equally to operational safety risk. Scenario-based testing, e.g., using optimisation algorithms to achieve high coverage, often overlooks this fact in its objectives, while mile-based testing naturally detects those bugs likely to manifest themselves in operation.
@@
iv) How does the fidelity [^3] While specific definitions of fidelity vary across studies, in general, fidelity refers to how closely synthetic data replicate real-world conditions, thereby supporting reliable simulation-based testing of AVs. of simulated scenarios affect safety estimates? Given that scenario-based testing focuses on high-risk conditions, synthetic data generated by simulation is typically employed to execute test scenarios. However, the Sim2Real gap introduces an additional layer of uncertainty, implying that safety claims based on synthetic scenario testing evidence may not fully capture real-world performance [10]. Despite great efforts having been made to bridge the gap [21, 22, 23, 24], there remains a lack of rigorous methods to quantify simulation fidelity [23] and translate varying fidelity levels into measurable safety risk.
@@
In this paper, we begin by reviewing the state-of-the-art in statistical modelling for AV safety assessment. We observe that most existing work focuses on mile-based testing, while statistical foundations for scenario-based testing remain largely under-explored, despite its growing prominence in industry practice. Furthermore, we notice that many of the open questions in scenario-based testing, e.g., effectiveness in debugging and residual risk estimation, closely parallel longstanding challenges studied in the traditional software testing community. We present an initial model, as a proof-of-concept, to demonstrate the potential of applying statistical approaches to rigorously reason about scenario-based safety evidence, and highlight the potential for AV testing to benefit from software testing, with existing models potentially being reusable. Our goal is to lay a foundation that future work can build upon to develop rigorous safety arguments for AVs.
@@
  • Proof-of-concept models using scenario-based testing to assess AV safety and quantify the reduced operational risk.
  • An example measure of Risk Estimation Fidelity of synthetic data, with formal definitions and discussion of its possible use towards simulation-based safety arguments.
@@ II-A AV Testing
Mile-based testing evaluates AV safety by accumulating real/simulated miles within the ODD, aiming to sample scenarios as they naturally occur. Safety is inferred statistically by analysing failure rates (e.g., crashes, disengagements) per unit of operation. It assumes an unbiased estimate and exposure to diverse scenarios (thus it may offers some benefits similar to scenario-based testing).
@@ II-B Software Testing
Broadly, software testing falls into two categories: directed testing and operational testing; they are also often called debug testing and acceptance testing [18, 26] based on their typical purposes. The former aims to find and fix defects, while the latter assesses existing quality to build confidence for deployment. The key difference is whether test cases are sampled from the operational profile (OP). Debug testing typically ignores the OP, whereas acceptance testing depends on it [18]. However, research shows operational testing may outperform directed testing in bug detection, and some acceptance tests rely solely on directed methods. We show this using our proof-of-concept models later.
@@
An OP is a probability distribution over the input space of a software system, representing the relative frequencies with which different input sequences expected occur during actual operation.
@@
Partition-based testing partitions the input domain of a software system into distinct “subdomains/partition”. Test cases are then selected from each subdomain, ensuring that at least one is selected from each subdomain.
@@ II-C Mapping Concepts Between AV and Software Testing
Scenario-based testing can be viewed as a form of partition-based testing, where logical scenarios partition the ODD, and representative concrete scenarios are picked for testing in each partition to ensure coverage. Both aim to save tests by assuming that with clever partitioning, failures will cluster in certain subdomains, allowing testers to focus effort where it matters most.
Remark 4 (Mile-based testing is random testing).
Continuing Remark 1, since each concrete scenario arises randomly from an operational scenario distribution (i.e., the OP), mile-based testing resembles randomly selected test cases (with replacement) across the scenario parameter space.
@@
That said, there are exceptional cases that we learnt from studies of software testing, e.g., [18, 19, 8, 17, 26, 28, 29]: 1) Partition/scenario-based testing is often thought to be good for debugging, but there is no formal proof and it may be less effective than operational testing under certain, not especially unusual, conditions. 2) Partition/scenario-based testing can still be used for estimating reliability (i.e., residual safety risk), if distributional information is explicitly used within, and across, logical scenario partitions. 3) Random/mile-based testing can still discover bugs, more or less efficiently, in terms of improving safety, than partition/scenario testing. All of these points are context-dependent and cannot be answered universally. This highlights the need for a theoretical model with statistical foundation capable of analysing different AV testing methods across conditions.
@@ III Statistical Modelling for AV Safety
[7, 36] pfm Bernoulli CBI Failure-free & rare failure illustrative data In addition to [6], prior knowledge may bring advantages, while avoiding optimistic biases.
@@
We can model the process of failure occurrence as occurring in a sequence of discrete trials (a discrete time model), so that the parameter of interest is a probability of failure per trip, scenario, etc., or in response to continuous exposure (e.g. number of miles driven or time driving), and in this continuous time model the parameter of interest is a rate of occurrence. The choice between discrete and continuous models determines the underlying stochastic process (e.g., Bernoulli for probabilities or Poisson for rates) and informs whether a discrete or continuous model is more appropriate. The metric also guides how data should be collected in practice, and whether the data realistically supports the assumptions of the chosen stochastic process.
@@
Various standards require “frequentist” estimates of confidence in bounds on the metric of interest. These may yield less confidence than desired, and incorrect estimates of the metric when too few data are available. Bayesian methods on the other hand can take into account prior knowledge about the AV, sometimes giving sufficient confidence despite limited data; but the reliability of the estimates depend on proper representation of the actual prior knowledge [43].
@@
SRGMs help track and predict failures over time, but they cannot be trusted a priori. They are especially problematic for highly reliable software and, even when generally accurate for a system, can occasionally produce seriously incorrect predictions.
@@ IV-A The Metric, Stochastic Failure Processes, and Estimators
As stated in Remark 6, the choice of metric is the first step in statistical inference. As a first attempt, we align with existing metrics and define the probability of failure per randomly selected scenario (drawn from the operational scenario distribution over the ODD), abbreviated as pfs:
@@
Similar to the “frequentist” interpretation of pfd in [44], pfs can be regarded as a limiting relative frequency of concrete scenarios for which the AV fails in an infinite sequence of independently selected concrete scenarios from the OP. Note, as stated in footnote 4, the “failure” in pfs is a general term that can be replaced by any safety event of the assessor’s interest, e.g., crash and near-miss.
@@ IV-A1 Scenario-based testing conducted in a mile-based/random testing way
likelihood function determines which statistical inference methods can be used to estimate $\theta$. These include those methods listed in Table I and the choice depends on the observed data (e.g., whether they include failures), prior knowledge available, and desired posterior estimation forms (e.g., mean or confidence bounds of pfs)
@@ IV-A2 Scenario-based testing conducted in a partition-based testing way
Let the scenario parameter space $D$ be partitioned into subdomains $D_{1},\dots,D_{n}$, i.e., $n$ logical scenarios. Then, by the law of total probability
@@
Thus, to estimate $\theta$, we need to know the conditional pfs for each logical scenario $\theta_{i}$ and the OP. Normally, inside a given logical scenario $D_{i}$, a new and (thought-to-be) more efficient distribution is used to generate concrete test scenarios (e.g., by an optimisation algorithm) rather than the (local) OP. Usually, the estimator using importance sampling can be used where the proposal distribution is either known or implicitly approximated [30, 31, 32, 33].
@@ IV-B Under Which Conditions Will Scenario-based Testing Deliver Better Safety than Mile-based Testing?
We continue to investigate the above question by retrofitting the model from [18] for AV testing. To model repair behaviour after detecting failures, [18] introduces the concept of “failure region ($F$)”—a set of inputs in $D$ that cause failure: a fix removes a specific failure region and does not introduce new failure regions [^6] The original paper [18] discusses the validity of some assumptions and model extensions to relax them, e.g., “partial fix”. For simplicity, we omit these discussions here.. While [18] considers the more general case of multiple (disjoint) failure regions, we only present the simplified single failure region case to better illustrate the conceptual mapping between AV and software testing, and the potential for reusing existing research insights.
Consider an AV under test whose pfs, $\theta=q$, is due to a single failure region. After mile-based testing that included $t$ random concrete scenarios (cf. Remarks 1 & 4), the failure region will have been detected, and fixed, or not. By the formula of total probability, the resulting expected pfs is as below, where the “00” reflects the fact that after fixing, no bugs remain.
@@
Consider the failure region is a strict subset of a single subdomain $k$, $F\subset D_{k}$, and with some points in $D_{k}$ not being failure points, $D_{k}\not\subset F$, and no other subdomains overlapping with $F$: $F\cap D_{i}=\emptyset,\forall i\neq k$. Additionally, we assume that within $D_{k}$, both scenario-based and mile-based testing (on average) have an equal probability of encountering $F$. That is, the detection probability $d_{k}$ is just the fraction of the operational distribution scenarios in $D_{k}$ that encounter $F$: $d_{k}=\frac{\sum_{x\in F}Op(x)}{\sum_{x\in D_{k}}Op(x)},$ For scenario-based testing, assume test budgets are distributed equally across $n$ subdomains, so $t_{k}=\frac{t}{n}$, where $t$ is the total number of tests.
  • Situations when scenario-based testing is superior: This case occurs when the failure-prone subdomain is relatively rare in real operation, resulting in fewer tests under mile-based testing compared to scenario-based testing. Mathematically, this is expressed as $t\sum_{x\in D_{k}}Op(x)\ll t_{k}$. After substituting $t_{k}=\frac{t}{n}$, we will have:
@@
    where the last term $q(1-q)^{t}$ is the pfs. The two approximations in Eq. (12) require that $d_{k}$ and $q$ are small, using $(1+x)^{y}\approx 1+yx$ for small $x$.
@@
  • Situations when mile-based testing is superior: A similar argument leads to the opposite conclusion when mile-based testing heavily samples $D_{k}$ (many mile-based test scenarios fall within $D_{k}$). If there are many other subdomains, scenario-based testing ends up “wasting” most of its effort on them (still assuming $t_{k}=t/n$), rather than the “failure prone” subdomain. Mathematically, mile-based testing provides more tests, $t\sum_{x\in D_{k}}Op(x)\gg t_{k}$, substituting $t_{k}=\frac{t}{n}$, we have: